OpenStudy (anonymous):

solve the system of equations 4x+7y=47 5x-4y=-5

jimthompson5910 (jim_thompson5910):

hint: multiply both sides of equation 1 by 4 and both sides of equation 2 by 7 after all this, add the equations

OpenStudy (anonymous):

wait what ._.

jimthompson5910 (jim_thompson5910):

4x+7y=47 5x-4y=-5 4*(4x+7y)=4*47 7*(5x-4y)=7*(-5) 16x+28y=188 35x-28y=-35

jimthompson5910 (jim_thompson5910):

now add the two equations 16x+28y=188 35x-28y=-35 ------------- 51x+0y = 153 51x = 153 x = ??

OpenStudy (anonymous):

10.2.?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

51x = 153 x = 153/51 x = 3

jimthompson5910 (jim_thompson5910):

use that to find y

OpenStudy (anonymous):

(5,3)

OpenStudy (anonymous):

(3,5)